# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ Appendix B Auxiliary Lemmas
For the first inequality, since for any matrix $A$ such that $\lVert A\rVert\leq q<1$ we have $(I-A)^{-1}=\sum_{n=0}^{\infty}A^{n}$ and hence $\lVert I-A\rVert\leq\sum_{n=0}^{+\infty}\lVert A\rVert^{n}\leq\sum_{n=0}^{+\infty}q^{n}=1/(1-q)$. Thus, if we let $\mathcal{A}=D_{\Phi,1}(u,\lambda)$ we have that
$\lVert(I-\mathcal{A})^{-1}\rVert_{\sup}=\sup_{A\in\mathcal{A}}\lVert(I-A)^{-1}\rVert\leq\frac{1}{1-q}.$
The second inequality holds since $D_{w}^{\operatorname{imp}}(\lambda)\subset D_{w}^{\operatorname{fix}}(\lambda)$. For the last inequality note that, from the definition of $D_{\Phi}^{\operatorname{fix}}$, we have that
$D_{\Phi}^{\operatorname{fix}}(u,\lambda)=D_{\Phi,1}(u,\lambda)D_{w}^{\operatorname{fix}}(\lambda)+D_{\Phi,2}(u,\lambda)$
and hence, from the properties of the excess in Lemma B.1 and from the fact that for every $v\in\mathbb{R}^{d}$ $\lVert D_{\Phi,1}(v,\lambda)\rVert_{\sup}\leq q<1$ we have
$\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}$ $\leq\lVert D_{\Phi,1}(u,\lambda)\rVert_{\sup}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+\lVert D_{\Phi,2}(u,\lambda)\rVert_{\sup}$
$\leq q\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+\lVert D_{\Phi,2}(u,\lambda)\rVert_{\sup}$
@@
We also present the following lemma, useful for the stochastic analysis.
Lemma B.4.
Under 3.1(ii) and 3.1, let $B_{\lambda}:=\lVert D_{\Phi,2}^{s}(w(\lambda),\lambda)\rVert_{\sup}$ and $C_{\lambda}$ be defined as in (12). Then, for every $(u,\lambda)\in\mathbb{R}^{p}\times\Lambda$, and also, with a slight abuse of notation set
$D_{\Phi}^{\operatorname{fix}}(u,\lambda)$ $:=D_{\Phi}^{\operatorname{fix}}(u,\lambda)D_{\Phi,1}(u,\lambda)+D_{\Phi,2}(u,\lambda)$
$D_{\Phi}^{\operatorname{imp}}(u,\lambda)$ $:=(I-D_{\Phi,1}(u,\lambda))^{-1}D_{\Phi,2}(u,\lambda).$
If we set $w=w(\lambda)$, We have
$\operatorname{\it{e}}(D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi}^{\operatorname{fix}}(w,\lambda))$ $\leq\frac{(B_{\lambda}+1)}{(1-q)^{2}}C_{\lambda}(u)\lVert u-w\rVert.$
The same holds also for $\operatorname{\it{e}}(D_{\Phi}^{\operatorname{imp}}(u,\lambda),D_{\Phi}^{\operatorname{imp}}(w,\lambda))$.
Proof .
Set, for the sake of brevity $w=w(\lambda)$.
Then, by definition we have
$D_{\Phi}^{\operatorname{fix}}(u,\lambda)=D_{\Phi,1}^{s}(u,\lambda)D_{\Phi}^{\operatorname{fix}}(u,\lambda)+D_{\Phi,2}^{s}(u,\lambda)\quad\text{and}\quad D_{\Phi}^{\operatorname{fix}}(w,\lambda)=D_{\Phi,1}^{s}(w,\lambda)D_{\Phi}^{\operatorname{fix}}(w,\lambda)+D_{\Phi,2}^{s}(w,\lambda).$
Thus, from the properties of the excess in Lemma B.1 we have
$\operatorname{\it{e}}(D_{\Phi}^{\operatorname{fix}}(u,\lambda),$ $D_{\Phi}^{\operatorname{fix}}(w,\lambda))$
$\leq\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda)D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi,1}^{s}(w,\lambda)D_{\Phi}^{\operatorname{fix}}(w,\lambda))+\operatorname{\it{e}}(D_{\Phi,2}^{s}(u,\lambda),D_{\Phi,2}^{s}(w,\lambda))$
$\leq\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda)D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi,1}^{s}(u,\lambda)D_{\Phi}^{\operatorname{fix}}(w,\lambda))$
$\quad+\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda)D_{\Phi}^{\operatorname{fix}}(w,\lambda),D_{\Phi,1}^{s}(w,\lambda)D_{\Phi}^{\operatorname{fix}}(w,\lambda))+\operatorname{\it{e}}(D_{\Phi,2}^{s}(u,\lambda),D_{\Phi,2}^{s}(w,\lambda))$
$\leq\lVert D_{\Phi,1}^{s}(u,\lambda)\rVert_{\sup}\operatorname{\it{e}}(D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi}^{\operatorname{fix}}(w,\lambda))$
$\quad+\lVert D_{\Phi}^{\operatorname{fix}}(w,\lambda)\rVert_{\sup}\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda),D_{\Phi,1}^{s}(w,\lambda))+\operatorname{\it{e}}(D_{\Phi,2}^{s}(u,\lambda),D_{\Phi,2}^{s}(w,\lambda))$
$\leq q\cdot\operatorname{\it{e}}(D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi}^{\operatorname{fix}}(w,\lambda))$
$\quad+\lVert D_{\Phi}^{\operatorname{fix}}(w,\lambda)\rVert_{\sup}\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda),D_{\Phi,1}^{s}(w,\lambda))+\operatorname{\it{e}}(D_{\Phi,2}^{s}(u,\lambda),D_{\Phi,2}^{s}(w,\lambda)).$ (23)
Hence
$\operatorname{\it{e}}(D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi}^{\operatorname{fix}}(w,\lambda))\leq\frac{\lVert D_{\Phi}^{\operatorname{fix}}(w,\lambda)\rVert_{\sup}}{1-q}\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda),D_{\Phi,1}^{s}(w,\lambda))+\frac{\operatorname{\it{e}}(D_{\Phi,2}^{s}(u,\lambda),D_{\Phi,2}^{s}(w,\lambda))}{1-q}.$
Now, by Lemma B.3 we have $\lVert D_{\Phi}^{\operatorname{fix}}(w,\lambda)\rVert_{\sup}\leq\lVert D_{\Phi,2}^{s}(w,\lambda)\rVert_{\sup}/(1-q)$ so that we get
$\operatorname{\it{e}}(D_{\Phi}^{\operatorname{fix}}(u,\lambda),D_{\Phi}^{\operatorname{fix}}(w,\lambda))$ $\leq\frac{B_{\lambda}}{(1-q)^{2}}\operatorname{\it{e}}(D_{\Phi,1}^{s}(u,\lambda),D_{\Phi,1}^{s}(w,\lambda))+\frac{\operatorname{\it{e}}(D_{\Phi,2}^{s}(u,\lambda),D_{\Phi,2}^{s}(w,\lambda))}{1-q}$
The statement follows from Lemma 3.2.
For the second part, let $\mathcal{A}_{i}=D_{\Phi,i}(u,\lambda)$ and $\mathcal{B}_{i}=D_{\Phi,i}(w,\lambda)$ for $i=1,2$ and recall that
$D_{\Phi}^{\operatorname{imp}}(u,\lambda)$ $=(I-D_{\Phi,1}(u,\lambda))^{-1}D_{\Phi,2}(u,\lambda)=(I-\mathcal{A}_{1})^{-1}\mathcal{A}_{2}$
$D_{\Phi}^{\operatorname{imp}}(w,\lambda)$ $=(I-D_{\Phi,1}(w,\lambda))^{-1}D_{\Phi,2}(w,\lambda)=(I-\mathcal{B}_{1})^{-1}\mathcal{B}_{2}.$
Then, using the properties in Lemma B.1(i)-(iii)-(v)-(vi), we get
$\operatorname{\it{e}}(D_{\Phi}^{\operatorname{imp}}(u,\lambda),$ $D_{\Phi}^{\operatorname{imp}}(w,\lambda))$
$\leq\operatorname{\it{e}}((I-\mathcal{A}_{1})^{-1}\mathcal{A}_{2},(I-\mathcal{A}_{1})^{-1}\mathcal{B}_{2})+\operatorname{\it{e}}((I-\mathcal{A}_{1})^{-1}\mathcal{B}_{2},(I-\mathcal{B}_{1})^{-1}\mathcal{B}_{2})$
$\leq\lVert(I-\mathcal{A}_{1})^{-1}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A}_{2},\mathcal{B}_{2})+\lVert\mathcal{B}_{2}\rVert_{\sup}\operatorname{\it{e}}((I-\mathcal{A}_{1})^{-1},(I-\mathcal{B}_{1})^{-1})$
$\leq\lVert(I-\mathcal{A}_{1})^{-1}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A}_{2},\mathcal{B}_{2})+\lVert\mathcal{B}_{2}\rVert_{\sup}\lVert(I-\mathcal{A}_{1})^{-1}\rVert_{\sup}\lVert(I-\mathcal{B}_{1})^{-1}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A}_{1},\mathcal{B}_{1})$
$\leq\frac{1}{1-q}\operatorname{\it{e}}(D_{\Phi,2}(u,\lambda),D_{\Phi,2}(w,\lambda))+\frac{\lVert D_{\Phi,2}(w,\lambda)\rVert_{\sup}}{(1-q)^{2}}\operatorname{\it{e}}(D_{\Phi,1}(u,\lambda),D_{\Phi,1}(w,\lambda))$
The statement follows again from Lemma 3.2. ∎
@@ Appendix C Iterative and Approximate Implicit Differentiation
Let $\lambda\in\Lambda$ and $t\in\mathbb{N}$, $t\geq 1$. We recall that
$D_{w_{t}}(\lambda)$ $=D_{\Phi,1}(w_{t-1}(\lambda),\lambda)D_{w_{t-1}}(\lambda)+D_{\Phi,2}(w_{t-1}(\lambda),\lambda),$
$D_{w}^{\operatorname{fix}}(\lambda)$ $=D_{\Phi,1}(w(\lambda),\lambda)D_{w}^{\operatorname{fix}}(\lambda)+D_{\Phi,2}(w(\lambda),\lambda).$
For simplicity, we let $b_{\lambda,t}=\big{(}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+1\big{)}C_{\lambda}(w_{t}(\lambda))$ and
$\mathcal{A}_{1,t}=D_{\Phi,1}(w_{t}(\lambda),\lambda),\ \mathcal{A}_{2,t}=D_{\Phi,2}(w_{t}(\lambda),\lambda),\ \mathcal{B}_{1}=D_{\Phi,1}(w(\lambda),\lambda),\ \mathcal{B}_{2}=D_{\Phi,2}(w(\lambda),\lambda).$
We also note that from the definition $E_{\lambda}(w_{t})=\mathds{1}\{\lVert w_{t}(\lambda)-w(\lambda)\rVert>R_{\lambda}\}$ and hence
$C_{\lambda}(w_{t})=L+\frac{M_{\lambda}}{R_{\lambda}}E_{\lambda}(w_{t})$
ITD (15): Let $\Delta^{\prime}_{t}:=\operatorname{\it{e}}(D_{w_{t}}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$. Using the properties of the excess in Lemma B.1 we have
$\Delta^{\prime}_{t}$ $=\operatorname{\it{e}}(\mathcal{A}_{1,t-1}D_{w_{t-1}}(\lambda)+\mathcal{A}_{2,t-1},\mathcal{B}_{1}D_{w}^{\operatorname{fix}}(\lambda)+\mathcal{B}_{2})$
$\leq\operatorname{\it{e}}(\mathcal{A}_{1,t-1}D_{w_{t-1}}(\lambda),\mathcal{A}_{1,t-1}D_{w}^{\operatorname{fix}}(\lambda)+\operatorname{\it{e}}(\mathcal{A}_{1,t-1}D_{w}^{\operatorname{fix}}(\lambda),\mathcal{B}_{1}D_{w}^{\operatorname{fix}}(\lambda))$
$\quad+\operatorname{\it{e}}(\mathcal{A}_{2,t-1},\mathcal{B}_{2})$
$\leq\lVert\mathcal{A}_{1,t-1}\rVert_{\sup}\Delta^{\prime}_{t-1}+\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A}_{1,t-1},\mathcal{B}_{1})+\operatorname{\it{e}}(\mathcal{A}_{2,t-1},\mathcal{B}_{2})$
$\leq q\Delta^{\prime}_{t-1}+b_{\lambda,t-1}\Delta_{t-1}$
where for the last inequality we used the contraction assumption and Lemma 3.2. By unrolling the recursive inequality and using the inequality $\Delta_{i}\leq q^{i}\Delta_{0}$ we obtain
@@
$\leq q^{t}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+\Delta_{0}q^{t-1}t(\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+1)(L+M_{\lambda}R^{-1}_{\lambda}\bar{E}_{\lambda}(t)),$
where, in the last inequality, we used $\Delta^{\prime}_{0}\leq\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}$ and the definitions of $\bar{E}_{\lambda}(t)$ and $C_{\lambda}(w_{t}(\lambda))$. Applying Lemmas B.3 and B.4, factoring out $t$ and using the definition of $B_{\lambda}$ gives the final result.
@@
$D_{w_{t}}^{k}(\lambda)=D_{\Phi,1}(w_{t}(\lambda),\lambda)D_{w_{t}}^{k-1}(\lambda)+D_{\Phi,2}(w_{t}(\lambda),\lambda).$
Set $\Delta^{\prime}_{k}:=\operatorname{\it{e}}(D_{w_{t}}^{k}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$. Then using again Lemma B.1 we have
$\Delta^{\prime}_{k}$ $=\operatorname{\it{e}}(\mathcal{A}_{1,t}D_{w_{t}}^{k-1}(\lambda)+\mathcal{A}_{2,t},\mathcal{B}_{1}D_{w}^{\operatorname{fix}}(\lambda)+\mathcal{B}_{2})$
$\leq\operatorname{\it{e}}(\mathcal{A}_{1,t}D_{w_{t}}^{k-1}(\lambda),\mathcal{A}_{1,t}D_{w}^{\operatorname{fix}}(\lambda))+\operatorname{\it{e}}(\mathcal{A}_{1,t}D_{w}^{\operatorname{fix}}(\lambda),\mathcal{B}_{1}D_{w}^{\operatorname{fix}}(\lambda))+\operatorname{\it{e}}(\mathcal{A}_{2,t},\mathcal{B}_{2})$
$\leq\lVert A_{1,t}\rVert_{\sup}\Delta^{\prime}_{k-1}+\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A}_{1,t},\mathcal{B}_{1})+\operatorname{\it{e}}(\mathcal{A}_{2,t},\mathcal{B}_{2})$
$\leq q\Delta^{\prime}_{k-1}+b_{\lambda,t}\Delta_{t}$
where for the last inequality we used Assumption 3.1(i) and Lemma 3.2. By unrolling the inequality recursion we obtain
$\Delta^{\prime}_{k}\leq q^{k}\Delta^{\prime}_{0}+b_{\lambda,t}\Delta_{t}\sum_{i=0}^{k-1}q^{i}=q^{k}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+b_{\lambda,t}\frac{1-q^{k}}{1-q}\Delta_{t}$
applying Lemmas B.3 and B.4 and using the definition of $b_{\lambda,t}$, $C_{\lambda}$ and $E_{\lambda}$ gives the final result.
For the final comment, if $w_{t}(\lambda)=\Phi(w_{t}(\lambda),\lambda)$, due to the contraction property of $\Phi$, $\Delta_{t}=q\Delta_{t-1}<\Delta_{t-1}$ and there exist $\tau_{\lambda}\in\{0,\dots,t^{\prime}_{\lambda}\}$ with $t^{\prime}_{\lambda}:=\lceil\log(\Delta_{0}/R_{\lambda})/\log(1/q)\rceil$, such that $\lVert w_{\tau_{\lambda}}(\lambda)-w(\lambda)\rVert\leq R_{\lambda}$, and if $\tau_{\lambda}\neq 0$, $\lVert w_{\tau_{\lambda}-1}(\lambda)-w(\lambda)\rVert>R_{\lambda}$. Thus, for every $i\in\mathbb{N}$ $E_{\lambda}(w_{i})=\mathds{1}\{i<\tau_{\lambda}\}$ and therefore for every $t$, $E_{\lambda}(w_{t})\leq\bar{E}_{\lambda}(t)\leq 1$.

### v2

@@ Appendix B Auxiliary Lemmas
As for the first inequality, we recall that for any matrix $A$ such that $\lVert A\rVert\leq q<1$, we have $(I-A)^{-1}=\sum_{n=0}^{\infty}A^{n}$ and hence $\lVert I-A\rVert\leq\sum_{n=0}^{+\infty}\lVert A\rVert^{n}\leq\sum_{n=0}^{+\infty}q^{n}=1/(1-q)$. Thus, if we let $\mathcal{A}=D_{\Phi}(u,\lambda)$ we have that
$\lVert(I-\mathcal{A}_{1})^{-1}\rVert_{\sup}=\sup_{A_{1}\in\mathcal{A}_{1}}\lVert(I-A_{1})^{-1}\rVert\leq\frac{1}{1-q}.$
The second inequality holds since $D_{w}^{\operatorname{imp}}(\lambda)\subset D_{w}^{\operatorname{fix}}(\lambda)$. For the last inequality we note that if we let $\mathcal{B}=D_{\Phi}(w(\lambda),\lambda)$, it follows from the definition of $D_{w}^{\operatorname{fix}}$ that
$D_{w}^{\operatorname{fix}}(\lambda)=\mathcal{B}(D_{w}^{\operatorname{fix}}(\lambda)).$
Thus, applying Lemma B.1(vii) and recalling that $\lVert D_{\Phi,1}(w(\lambda),\lambda)\rVert_{\sup}\leq q<1$ we have
$\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}$ $\leq\lVert D_{\Phi,1}(w(\lambda),\lambda)\rVert_{\sup}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}$
$\leq q\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}$
@@ Appendix C Iterative and Approximate Implicit Differentiation
Let $\lambda\in\Lambda$ and $t\in\mathbb{N}$, $t\geq 1$. For the sake of brevity, we set
$b_{\lambda,t}$ $=\big{(}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+1\big{)}C_{\lambda}(w_{t}(\lambda)),$
$\mathcal{A}_{t}$ $=D_{\Phi}(w_{t}(\lambda),\lambda),\quad\mathcal{A}_{t,1}=D_{\Phi,1}(w_{t}(\lambda),\lambda),\quad\mathcal{B}=D_{\Phi}(w(\lambda),\lambda),$
where $C_{\lambda}$ is defined in Lemma 3.2. We recall that
$D_{w_{t}}(\lambda)$ $=\mathcal{A}_{t-1}(D_{w_{t-1}}(\lambda)),\qquad D_{w}^{\operatorname{fix}}(\lambda)=\mathcal{B}(D_{w}^{\operatorname{fix}}(\lambda)).$
We also recall that $\delta_{\lambda}(t)=\mathds{1}\{\lVert w_{t}(\lambda)-w(\lambda)\rVert>R_{\lambda}\}\in\{0,1\}$ and hence
$C_{\lambda}(w_{t}(\lambda))=L+\frac{M_{\lambda}}{R_{\lambda}}\delta_{\lambda}(t).$
ITD (16): Let $\Delta^{\prime}_{t}:=\operatorname{\it{e}}(D_{w_{t}}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$. Using the properties in Lemma B.1(i)(vii) we have
$\Delta^{\prime}_{t}$ $=\operatorname{\it{e}}(\mathcal{A}_{t-1}(D_{w_{t-1}}(\lambda)),\mathcal{B}(D_{w}^{\operatorname{fix}}(\lambda)))$
$\leq\operatorname{\it{e}}(\mathcal{A}_{t-1}(D_{w_{t-1}}(\lambda)),\mathcal{A}_{t-1}(D_{w}^{\operatorname{fix}}(\lambda)))+\operatorname{\it{e}}(\mathcal{A}_{t-1}(D_{w}^{\operatorname{fix}}(\lambda)),\mathcal{B}(D_{w}^{\operatorname{fix}}(\lambda)))$
$\leq\lVert\mathcal{A}_{t-1,1}\rVert_{\sup}\Delta^{\prime}_{t-1}+(1+\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup})\operatorname{\it{e}}(\mathcal{A}_{t-1},\mathcal{B})$
$\leq q\Delta^{\prime}_{t-1}+b_{\lambda,t-1}\Delta_{t-1},$
where for the last inequality we used that for any $u\in\mathbb{R}^{d}$, $\lVert D_{\Phi,1}(u,\lambda)\rVert<q$ and Lemma 3.2. By unrolling the recursive inequality and using the inequality $\Delta_{i}\leq q^{i}\Delta_{0}$ we obtain
@@
$\leq q^{t}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+\Delta_{0}q^{t-1}t(\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+1)(L+M_{\lambda}R^{-1}_{\lambda}\bar{\delta}_{\lambda}(t)),$
where, in the last inequality, we used $\Delta^{\prime}_{0}\leq\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}$ and the definitions of $\bar{\delta}_{\lambda}(t)$ and $C_{\lambda}(w_{t}(\lambda))$. Applying Lemma B.3, factoring out $t$ and using the definition of $B_{\lambda}$ gives the final result.
@@
$D_{w_{t}}^{k}(\lambda)=\mathcal{A}_{t}(D_{w_{t}}^{k-1}(\lambda)).$
Set $\Delta^{\prime}_{k}:=\operatorname{\it{e}}(D_{w_{t}}^{k}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$. Then using again Lemma B.1(i)(vii) we have
$\Delta^{\prime}_{k}$ $=\operatorname{\it{e}}(\mathcal{A}_{t}(D_{w_{t}}^{k-1}(\lambda)),\mathcal{B}(D_{w}^{\operatorname{fix}}(\lambda)))$
$\leq\operatorname{\it{e}}(\mathcal{A}_{t}(D_{w_{t}}^{k-1}(\lambda)),\mathcal{A}_{t}(D_{w}^{\operatorname{fix}}(\lambda)))+\operatorname{\it{e}}(\mathcal{A}_{t}(D_{w}^{\operatorname{fix}}(\lambda)),\mathcal{B}(D_{w}^{\operatorname{fix}}(\lambda)))$
$\leq\lVert\mathcal{A}_{t,1}\rVert_{\sup}\Delta^{\prime}_{k-1}+(1+\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup})\operatorname{\it{e}}(\mathcal{A}_{t},\mathcal{B})$
$\leq q\Delta^{\prime}_{k-1}+b_{\lambda,t}\Delta_{t},$
where for the last inequality we used Assumption 3.1(ii) and Lemma 3.2. By unrolling the inequality recursion we obtain
$\Delta^{\prime}_{k}\leq q^{k}\Delta^{\prime}_{0}+b_{\lambda,t}\Delta_{t}\sum_{i=0}^{k-1}q^{i}=q^{k}\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}+b_{\lambda,t}\frac{1-q^{k}}{1-q}\Delta_{t}.$
Applying Lemma B.3 and using the definition of $b_{\lambda,t}$, $C_{\lambda}$ and $\delta_{\lambda}$ gives the final result.
For the final comment, if $w_{t}(\lambda)=\Phi(w_{t-1}(\lambda),\lambda)$, due to the contraction property of $\Phi$, $\Delta_{t}=q\Delta_{t-1}<\Delta_{t-1}$ and there exist $\tau_{\lambda}\in\{0,\dots,t^{\prime}_{\lambda}\}$ with $t^{\prime}_{\lambda}:=\lceil\log(\Delta_{0}/R_{\lambda})/\log(1/q)\rceil$, such that $\lVert w_{\tau_{\lambda}}(\lambda)-w(\lambda)\rVert\leq R_{\lambda}$, and if $\tau_{\lambda}\neq 0$, $\lVert w_{\tau_{\lambda}-1}(\lambda)-w(\lambda)\rVert>R_{\lambda}$. Thus, for every $i\in\mathbb{N}$ $\delta_{\lambda}(w_{i})=\mathds{1}\{i<\tau_{\lambda}\}$ and therefore for every $t$, $\delta_{\lambda}(w_{t})\leq\bar{\delta}_{\lambda}(t)\leq 1$. ∎